OpenStudy (yanasidlinskiy):

The width w of a rectangular swimming pool is 7x^2. The area A of the pool is 7x^3 – 42x^2. What is an expression for the length of the pool?

OpenStudy (yanasidlinskiy):

@hartnn Can you help me?

OpenStudy (paki):

width(w) = 7x^2 Area(A) = 7x^3 – 42x^2 = 7x^2(x-6) Since... Area = width*length Area/width = length or [7x^2(x-6)]/[7x^2] = length (x-6) = length x-6 = length (this is what they're looking for)... @YanaSidlinskiy